OpenStudy (math2400):

can someone just tell me if these calc answers are right:

OpenStudy (math2400):

for the first one i did this: \[\frac{ 1 }{ 5-0 }\int\limits_{0}^{5} (60x ^{3}-650x ^{2}+1200x+4500) dx\]

jimthompson5910 (jim_thompson5910):

The first one is correct. I'm getting 11875/3 = 3958.333333

jimthompson5910 (jim_thompson5910):

the second one is not correct

OpenStudy (math2400):

perfect same here but I'm not sure if i did it right. Did u use the integral i provided or did u come up with it on your own and then integrate? Now for the second one i used this but i'm honestly not sure

jimthompson5910 (jim_thompson5910):

I integrated what you wrote

OpenStudy (math2400):

oh but i'm not sure if my integral is right lol

jimthompson5910 (jim_thompson5910):

your integral for the first one is correct

jimthompson5910 (jim_thompson5910):

how did you get the answer to the second one?

OpenStudy (math2400):

i used this integral: \[\int\limits_{-1}^{1} (2-x ^{2})^{2}-(9x ^{2})^{2} dx\]

jimthompson5910 (jim_thompson5910):

you are supposed to have \[\Large \int_{-1}^{1} \left[(2-x^2)-(x^2)\right]^2dx\] notice how I subtracted the functions first, and then squared the difference

jimthompson5910 (jim_thompson5910):

and idk where you got the 9 from

OpenStudy (math2400):

my bad the 9 was an accident i was trying to the the parenthesis but i didn't hold shift haha

OpenStudy (math2400):

and ok that's because it's a volume right ?

jimthompson5910 (jim_thompson5910):

ah gotcha lol

jimthompson5910 (jim_thompson5910):

the integrand \[\Large \left[(2-x^2)-(x^2)\right]^2\] represents the area of any cross section. When you combine those areas (like stack them up) by integrating, they form a 3D solid volume

OpenStudy (math2400):

ok i see. So the answer is 4.26667

jimthompson5910 (jim_thompson5910):

write that as a fraction

OpenStudy (math2400):

64/15 sorry my bad i took so long. Phone call!

jimthompson5910 (jim_thompson5910):

that's perfectly fine

jimthompson5910 (jim_thompson5910):

64/15 is correct

OpenStudy (math2400):

and i'm assuming the last one is correct right? because u square it after ?

jimthompson5910 (jim_thompson5910):

still checking that one

jimthompson5910 (jim_thompson5910):

I'm guessing there's a 4th answer choice that is cut off? Or no?

OpenStudy (math2400):

ya the forth answer is "none of the above"

jimthompson5910 (jim_thompson5910):

ok thanks

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):